Mathematics 241 – Multivariable Calculus Quiz 8 – November 16, 2007 (the final quiz of the semester!)

Your Name:

(5) A) Evaluate ∫ ∫_{R} x − 2 dA where R is the rectangle R = [0, 1] × [1, 2] in R^{2}.

(5) B) Sketch the region of integration for the integral ∫_{0}^{2} ∫_{x}^{2x} f(x, y) dy dx, and interchange the order of integration (i.e. write as one or more integrals ∫ ∫ f(x, y) dx dy).